Now suppose youhave two stars, with one them 2times more luminous than the other. From Earth you see that they both appear to have exactly the same apparent brightness or intensity. How much farther away is the bright star compared to the fainter one?
Answer #1

as the intensity of the light reaching

Intensity = Power/(4pi * d^2)

as bright star is 2 times brighter

the power will be 2 times

P1/d1^2 = P2/d2^2

>> P1/P2 = 2 = (d1/d2)^2

d1/d2 = sqrt(2)

the bright star is sqrt(2) times or 1.414 times the fainter star